### v3

# Computing Efficiently in QLDPC Codes

Abstract
It is the prevailing belief that quantum error correcting techniques will be required to build a utility-scale quantum computer able to perform computations that are out of reach of classical computers. The quantum error correcting codes that have been most extensively studied and therefore highly optimized, surface codes, are extremely resource intensive in terms of the number of physical qubits needed. A promising alternative, quantum low-density parity check (QLDPC) codes, has been proposed more recently. These codes are much less resource intensive, requiring significantly fewer physical qubits per logical qubit than practical surface code implementations. A successful application of QLDPC codes would therefore drastically reduce the timeline to reaching quantum computers that can run algorithms with exponential speedups like Shor’s algorithm and Quantum Phase Estimation (QPE). However to date QLDPC codes have been predominantly studied in the context of quantum memories; there has been no known method for implementing arbitrary logical Clifford operators in a QLDPC code proven efficient in terms of circuit depth. In combination with known methods for implementing $T$ gates, an efficient implementation of the Clifford group unlocks resource-efficient universal quantum computation. In this paper, we introduce a new family of QLDPC codes that enable efficient compilation of the full Clifford group via transversal operations. Our construction executes any $m$-qubit Clifford operation in at most $O(m)$ syndrome extraction rounds, significantly surpassing state-of-the-art lattice surgery methods. We run circuit-level simulations of depth-126 logical circuits to show that logical operations in our QLDPC codes attains near-memory performance. These results demonstrate that QLDPC codes are a viable means to reduce the resources required to implement all logical quantum algorithms, thereby unlocking a reduced timeline to commercially valuable quantum computing.
I Introduction
Quantum computers are poised to deliver the next major evolution in computational technology, with known applications in several high-impact sectors, including drug discovery, materials design, and cryptanalysis. However, by their nature, quantum technologies are highly susceptible to noise and are inherently incompatible with classical error correction techniques, leading to the development of unique quantum-specific error correction solutions. In quantum error correction (QEC), physical redundancy is introduced so that errors can be detected and corrected, ideally without harming the encoded information. QEC researchers have spent decades optimizing so-called “planar” QEC codes, such as the surface code. These codes have many positive attributes, including nearest-neighbour connectivity for syndrome extraction, competitive error thresholds, and a high degree of symmetry. However, the physical resource requirements of the surface code are of surface code error correction are so onerous that experts have assessed that with these codes quantum technologies capable of breaking RSA-2048 (a benchmark often used to assess commercial utility) are likely to only arrive decades in the future [1]. Until efficient QEC is unlocked at scale, commercial quantum applications will not deliver material value to society.
Tantalizingly, there are other QEC codes –known as quantum low-density parity check (QLDPC) codes  [^1] Note that throughout this paper we follow the convention of the field and implicitly exclude surface codes from consideration when discussing QLDPC codes, instead focusing on codes with higher encoding rate that meet the LDPC criteria.– that do not suffer from such high physical overheads. These codes have many of the same positive traits that planar codes do, leading to an explosion of interest in them over the past few years [3, 4, 5, 6, 7, 8, 9, Brown.2022, 11]. However, despite significant work in this direction [9, 6, 12], it is not presently known how to compile depth-efficient logical operations in QLDPC codes. If provably efficient universal logical gate sets were to be found for QLDPC codes, the timeline for the availability of commercially relevant quantum computers could be brought in by years if not decades, owing to the reduced cost in physical qubit count.
In this paper, we propose QLDPC codes, Subsystem HYpergraph Product Simplex (SHYPS) codes (see Fig. 1), designed from the bottom up with logical gate implementation as the core consideration. We construct highly symmetric codes capable of implementing immense numbers of logical gates transversally [13], facilitating arbitrary logic with asymptotic circuit depths matching those for unencoded logic. This approach is generally compatible with single-shot error correction, providing benefits in reduced qubit-overhead, reduced decoding complexity, and faster logical clock speeds. We demonstrate this compatibility by performing the most advanced circuit-level logical simulation to date, demonstrating near-memory performance for a compiled circuit derived from a randomly sampled logical Clifford circuit on 18 logical qubits.
Figure 1: Gauge generator connectivity for $r=3$ SHYPS code. Black circles denote data qubits. Orange (blue) squares indicate auxiliary qubits used to measure $X$ ($Z$) gauge generators. Each gauge generator has weight 3 and is measured using three CNOT gates, depicted as edges connecting the corresponding data qubits to a single auxiliary qubit. Two representative gauge generators are illustrated: one $X$-type and one $Z$-type. The full set of gauge generators can be obtained by translating this pattern under periodic boundary conditions.
II Logic in QLDPC Codes
The Clifford group is of particular interest for quantum computation as it provides a universal gate set when combined with any single non-Clifford operator. As there are known methods for implementing non-Clifford operations (e.g., $T$ gates) fault-tolerantly using state injection [14], an efficient fault-tolerant implementation of the Clifford group for a given code is sufficient to unlock universal computation in that code.
To date, QLDPC codes have been predominantly studied in the context of quantum memories. Various techniques based on generalized lattice surgery [10, 7] have been proposed to rework logical Clifford implementations in planar codes for application in QLDPC codes. However, these methods often introduce substantial overheads in both time and space: operations are not guaranteed to be parallelizable, large auxiliary patches are required, and any single-shot properties [11] of the QLDPC code can be compromised, as lattice surgery does not inherently support them.
In the generalized lattice surgery framework, computation is performed via joint logical operator measurements. State-of-the-art schemes for QLDPC surgery [6] measure single weight-$d$ logical operators with overheads of $O(d)$ in both space and time for a single logical cycle, where $d$ is the code distance. A constant number of such measurements may then be combined to implement elementary Clifford gates such as CNOT, Hadamard and Phase. It may be possible that compiling methods will be developed to optimize depth-efficiency for these techniques. However, even assuming a high degree of parallelization commensurate with planar codes –which may or may not be possible in general– direct compilation with these elementary gates to implement a worst-case $m$-qubit Clifford operator requires $O(m)$ logical cycles, leading to a total depth of $O(md)$.
Recent research [8] suggests that logical Clifford execution based on transversal gates may avoid the $O(d)$ overhead for a single logical cycle, allowing syndrome information to accrue simultaneously with logical Clifford execution. This indicates that a QLDPC code family with sufficiently many transversal gates that also exhibits single-shot capabilities might execute arbitrary Clifford logic in lower depths with efficient decoding strategies. We generate a new family of highly symmetric QLDPC codes –Subsystem HYpergraph Product Simplex (SHYPS) codes– by combining the subsystem hypergraph product (SHP) code construction [15] with classical simplex codes [book:MW, Ch. 1.9], and show that for these codes such low-depth logical Clifford implementations are indeed possible. The SHYPS code family permits implementations of elementary Clifford gates in $O(1)$ time with typically zero space overhead (see Table 1). However, the chief figure of merit we analyze is the depth of a worst-case encoded Clifford.
Using new compilation techniques, we achieve a depth-efficient implementation of a worst-case $m$-qubit Clifford in only $4m(1+o(1))$ logical cycles, each consisting of a single depth-1 circuit followed by a depth-6 syndrome extraction round. This is comparable to state-of-the-art depth-efficiency, $2m+O(log^{2}(m))$, for a purely unitary implementation of Cliffords acting on unencoded qubits [17]. The compilation strategy employed to achieve this low-depth implementation constructs circuits not as a sequence of elementary one- and two-qubit logical operations, but rather by leveraging many-qubit logical operations that occur naturally via low physical depth logical generators. Circuit-level simulations of these compiled circuits show near-memory performance (see Fig. 3), demonstrating the feasibility of time-efficient logical execution in QLDPC codes.
III Compiling and Costs
When computing in an error correction setting, logical operations are interleaved with rounds of syndrome extraction  [^2] Syndrome extraction results inform error correction, and corrections are often tracked in software and folded into the operators applied later in the circuit. For the purposes of this discussion, we restrict attention to syndrome extraction as this is the piece that is costly in terms of operations applied to qubits.. The operators are typically drawn from a subset of the full set of logical operations, and we will refer to elements of this subset as logical generators. A sequence of logical generators combine to synthesize a logical operation that the circuit performs. To compute efficiently in a QEC code, a generating set of (ideally low-depth) logical generators is needed. The size of this generating set relative to the space of logical operations it generates, informs the worst-case number of generators needed to implement an arbitrary logical operation, and therefore the number of syndrome extraction steps that need to be interleaved between them. Since syndrome extraction is costly and lengthier computations are less desirable, the aim of compiling is to reduce the number of steps needed to implement a desired logical circuit. This reduction can be achieved by constructing codes which have more native low-depth logical generators. We restrict attention to the space of the $m$-qubit Clifford group, and examine the notion of efficient compilation in that context to quantify the number of logical generators needed to compute efficiently.
The size of the $m$-qubit Clifford group scales as $2^{2m^{2}+O(m)}$ [19] (see also Section IV.1), which tends to pose a problem for compiling with a number of logical generators that doesn’t also grow at least exponentially in $m$. Consider the case where we work with $\gamma$ fixed-depth logical generators of the $m$-qubit Clifford group. The set of all circuits composed of up to depth $D$ of these logical generators cannot produce more than $(\gamma+1)^{D}$ unique Clifford operators. Ignoring the action of the Pauli group (all Paulis can be pushed to the end and implemented via a depth-1 circuit in stabilizer codes), this means we need $D$ to be such that
$(\gamma+1)^{D}\geq 2^{2m^{2}+O(m)}$
to possibly produce every Clifford operator. Rearranging, this lower bounds the required depth to attain an arbitrary operation in the Clifford group:
$D\geq\frac{2m^{2}+m-1}{\log_{2}(\gamma+1)}=:D^{*}.$ (1)
We can estimate that the fraction of total Clifford operators achievable with all depths $D<D^{*}$ is $1/(\gamma+1)$. Since we generally consider cases where $\gamma$ is a monotonically increasing function of $m$, asymptotically we expect $1/(\gamma+1)\rightarrow 0$ so that almost all Cliffords require depth at least $D^{*}$.
Consider synthesizing $m$-qubit Clifford operators using depth-1 circuits of arbitrary one- and two-qubit Clifford gates. These circuits are fixed under conjugation by the Clifford subgroup $\left<S,H,\text{SWAP}\right$>, and a counting exercise demonstrates that there are $\gamma\leq e/\sqrt{\pi}\left(10m/e\right)^{m/2}-1$ nontrivial depth-1 circuits, up to the right action of this subgroup. Substituting this expression into Eq. 1 yields an asymptotic worst-case (and average-case) depth lower bound of $D\geq 4m/\log_{2}(10m/e)+O(1)$. While there are indeed compiling algorithms that achieve depth $O(m/\log_{2}m)$ asymptotically [20], the leading constants are impractically large [17]. Rather than accept these large overheads, researchers have derived alternative decompositions that achieve better depths in practice for realistic qubit counts with worst-case depths of $O(m)$ [17, 21, 22, 13].
We employ an alternative decomposition (see Supplementary Material LABEL:supmat:sec:SHSCompSummary), which focuses on minimizing the rounds of specific subsets of Clifford gates. Every Clifford operator can be written in the five-stage decomposition $\mathrm{DZ}-\mathrm{DX}-\mathrm{DZ}-\mathrm{DX}-\mathrm{DZ}(1)$  [^3] One can consider this decomposition for the inverse or the Hadamard-conjugated version of a Clifford operator to find a total of four related decomposition (see Supplementary Materials LABEL:supmat:sec:SHSCompSummary for details). This freedom in reordering allows compilers to leverage interaction with neighbouring T gates effectively. , where we have defined subsets of the Clifford group as follows:
  • $\mathrm{DZ}=\langle S,CZ\rangle$; $Z$-diagonal Cliffords,
  • $\mathrm{DX}$; $X$-diagonal Cliffords (Hadamard-rotated $\mathrm{DZ}$),
  • $\mathrm{DZ}(1)$; depth-1 $Z$-diagonal operations.
Combined with a few additional insights [17], any of the listed decompositions yield circuits with asymptotic depths of $2m+O(\log^{2}m)$, which are state of the art for realistic qubit counts. Compiling worst case logical circuits on $m$ patches of surface code with all-to-all connectivity for an arbitrary $m$-qubit logical Clifford will thus require roughly $2m$ logical cycles.
Having established via study of decompositions a worst-case depth for a logical Clifford in the surface code, we can now revisit Eq. 1 and consider the number of logical Clifford generators we require for a different error correcting code to achieve similar efficiency. Clearly, $\gamma\sim 2^{O(m)}$ is necessary to achieve a logical cycle depth of $O(m)$. However, note that because CSS codes always have transversal CNOT operations available as logical generators, when using large numbers of code blocks of a CSS code, we will always satisfy this requirement. This holds because the number of ways to pair up code blocks to apply cross-block operations scales exponentially in the number of code blocks (and hence the total number of logical qubits). Instead, we would like to capture the compiling behavior in code families at both the “few” code block scale and the “many” code block scale.
For $b$ code blocks of an $[n,k,d]$ code, the Clifford compiling ratio
[…]
captures compiling properties for an arbitrary number of code blocks  [^4] Note that the focus on worst case Cliffords is equivalently a focus on average case Cliffords. While we might be tempted to introduce a metric that looks more directly at the worst-case depth increase for a specific Clifford operator, it turns out that such a metric is not particularly useful outside this instance. In particular, while there are circuits that achieve low depths in an unencoded setting that require high depths in the encoded setting, there will in general also be logical circuits that require low physical depths to implement whose corresponding unencoded counterpart only has high depth implementations. Since the worst case Clifford depth comes from a purely combinatoric argument, these code-specific details become irrelevant, allowing for a straightforward comparison with real predictive power.. Note that, throughout this paper, we adopt the following notation for error-correcting codes to ensure clarity: $(n,k,d)$ for classical codes, $[n,k,d]$ for subsystem codes, and $[[n,k,d]]$ for stabilizer codes. Any code that achieves an $O(1)$ ratio is said to generate the logical Clifford group efficiently. A code family with parameters $[n(r),k(r),d(r)]$ for which every member generates the logical Clifford group efficiently is also said to have this feature. Any code family whose associated compiling is such that the Clifford compiling ratio scales with $k(r)$ is failing to keep pace with the surface code due to the depth of logical operations within a code block, whereas scaling with $b$ is associated with overheads for compiling between code blocks.
Motivated by the advantages of computing in a code that has many fault-tolerant logical generators of low depth, we introduce the SHYPS code family in Section V. This family has $\gamma=2^{O(k)}$ logical generators for a single code block, each implemented by a depth-1 physical circuit. Moreover, these logical generator implementations often require 0 additional qubits, rising to at most $n$ for in-block CNOT operators where a scheme involving an auxiliary code block is used  [^5] There exist methods to remove the need for this additional auxiliary block with the same asymptotic cost, but the actual circuit length tends to be larger in the low code block regime.
In addition to possessing a sufficient number of logical generators, the SHYPS codes achieve the desired $O(1)$ Clifford compiling ratio, with logical Clifford operators across $b$ blocks implemented fault-tolerantly in depth at most $4bk(1+o(1))$ (see Table 1). Crucially, the depth of Clifford operations compiled in our SHYPS code framework remains independent of the code distance, compared to state-of-the-art lattice surgery methods where the depth scales as $O(md)$. For a moderately sized code with distance $d=20$, an SHYPS-compiled CNOT gate would achieve an order-of-magnitude reduction in depth relative to the equivalent compiled using lattice surgery ($4$ vs. roughly $40$). This example highlights that –in addition to reducing qubit overheads relative to the surface code– an SHYPS-based quantum computer would provide substantially faster clock-speeds at the logical level. The efficient Clifford compiling and substantially lower qubit overheads of SHYPS codes result in a lower space-time volume for Clifford operators in SHYPS compared to surface codes with either lattice surgery or transversal gates.
Logical Gate Time cost Space cost
CNOT (cross-block) 4 0
CNOT (in-block) 4 $n$
$S$ (in-block) 6 0
CZCZ (cross/in-block) 4 0
$H$ (in-block) 8 0
Arbitrary $b$-block Clifford $4bk(1+o(1))$ 0
Table 1: Time and space costs for logical Clifford operations of $SHYPS(r)$ codes ($r\geq 4$) with parameters $[n,k]$. For $r=3$, the logical $S$ and $H$ gates have depths 99 and $11$, respectively. Here, time corresponds to the number of complete syndrome extraction rounds.
IV Clifford operators and automorphisms
IV.1 Symplectic representations
The Clifford group $\mathcal{C}_{n}$ is a collection of unitary operators that maps the Pauli group $\mathcal{P}_{n}$ to itself, under conjugation. For example, the two-qubit controlled-not operator $CNOT_{i,j}$, the single-qubit phase gate $S_{i}$, and the single-qubit Hadamard gate $H_{i}$, are all Clifford operators, and in fact these suffice to generate the full group. When considering logical gates of codes, it is convenient to utilise the well-known symplectic representation of Clifford gates: by definition $\mathcal{P}_{n}$ is a normal subgroup of $\mathcal{C}_{n}$ and the quotient $\mathcal{C}_{n}/\mathcal{P}_{n}$ is isomorphic to $Sp_{2n}(2)$, the group of $2n\times 2n$ binary symplectic matrices [Rengaswamy2018, Thm. 15]. Hence each Clifford operator is, up to Pauli, specified by a unique element $g\in Sp_{2n}(2)$. That this representation ignores Pauli factors is of no concern, as for stabilizer codes, any logical Pauli may be implemented transversally.
The following examples illustrate the symplectic representation of some common families of Clifford operators; note that by convention we assume that elements of $Sp_{2n}(2)$ act on row vectors from the right.
Example IV.1.
(CNOT circuits) The collection of CNOT circuits $\left<CNOT_{i,j}\,:\,1\leq i,j\leq n\right$> have symplectic representations
$\left\{\begin{bmatrix}C&0\\ 0&C^{-T}\end{bmatrix}\,:\,C\in GL_{n}(2)\right\},$
where $GL_{n}(2)$ is the group of invertible $n\times n$ binary matrices.
Example IV.2.
(Diagonal Clifford operators) The Clifford operators that act diagonally on the computational basis form an abelian group, generated by single-qubit phase gates $S_{i}$ and the two-qubit controlled-Z gate $CZ_{i,j}$. They are represented by symplectic matrices of the form
$\left\{\begin{bmatrix}I_{n}&B\\ 0&I_{n}\end{bmatrix}\,:\,B\in M_{n}(2),B^{T}=B\right\},$
where the diagonal and off-diagonal entries of the symmetric matrix $B$, determine the presence of $S$ and CZ gates, respectively.
IV.2 Code automorphisms
Code automorphisms are permutations of the physical qubits that preserve the codespace. They are a promising foundation for computing in QLDPC codes as they can provide nontrivial logical gates implementable by low-depth SWAP circuits, or simply relabelling physical qubits. Moreover, combining automorphisms with additional transversal Clifford gates can give greater access to fault-tolerant logical gate implementations [3, 4, 9].
Let $\mathcal{C}$ be an $[n,k,d]$ CSS code with $X$- and $Z$-type gauge generators determined by matrices $G_{X}\in\mathbb{F}_{2}^{r_{X}\times n}$ and $G_{Z}\in\mathbb{F}_{2}^{r_{Z}\times n}$, respectively. The (permutation) automorphism group ${\mathrm{Aut}}(\mathcal{C})$ is the collection of permutations $\pi\in S_{n}$ that preserve the gauge generators, and therefore the codespace. I.e., $\pi\in{\mathrm{Aut}}(\mathcal{C})$ if there exist $g_{\pi,X}\in GL_{r_{X}}(2)$ and $g_{\pi,Z}\in GL_{r_{Z}}(2)$ such that $g_{\pi,X}G_{X}=G_{X}\pi$ and $g_{\pi,Z}G_{Z}=G_{Z}\pi.$
The logical gate implemented by a given $\pi\in{\mathrm{Aut}}(\mathcal{C})$ is determined by its action on the code’s logical Pauli operators. In particular, as permutations preserve the $X/Z$-type of a Pauli operator, $\pi$ implements a logical CNOT circuit [Grassl2013, Thm. 2]. Furthermore, following [26], a larger set of fault-tolerant CNOT circuits across two copies of $\mathcal{C}$ may be derived by conjugating the target block of the standard transversal CNOT operator  [^6] The physical transversal CNOT operator implements logical transversal CNOT in any subsystem CSS code [shor1997, Sec. 5] by $\pi$ (see Supplementary Materials LABEL:supmat:fig:generalised-transversal-cnot).
The symplectic representation for this combined operator is given by
$\begin{bmatrix}I_{n}&\pi&\\ 0&I_{n}&&\\ &&I_{n}&0\\ &&\pi^{-1}&I_{n}\\ \end{bmatrix}\in Sp_{4n}(2),$ (2)
where we identify $\pi\in{\mathrm{Aut}}(\mathcal{C})$ with the permutation matrix in $GL_{n}(2)$ whose $(i,j)$-th entry is 1 if $i=\pi(j)$, and zero otherwise. Note that as conjugation by $\pi$ simply permutes the targets of the physical transversal CNOT, this is a fault-tolerant circuit of depth 1. Moreover, (2) implements a cross-block CNOT operator, with all controls in the first code block of $k$ logical qubits, and all targets in the latter code block.
More recently, code automorphisms have been generalized to include qubit permutations that exchange vectors in $G_{X}$ and $G_{Z}$. These so-called $ZX$-dualities lead to low-depth logical operator implementations involving qubit permutations and single qubit Hadamard gates [3]. Moreover, $ZX$-dualities allow for the construction of logical diagonal Clifford operators in the following manner: Let $\tau\in S_{n}$ be an involution ($\tau^{2}=1$) such that $G_{X}\tau=G_{Z}$, and suppose that $\pi\in{\mathrm{Aut}}(\mathcal{C})$ is such that $\pi\tau$ is also an involution. Then the physical diagonal Clifford operator given by
$\begin{bmatrix}I_{n}&\pi\tau\\ 0&I_{n}\end{bmatrix}\in Sp_{2n}(2),$ (3)
is a depth-1 circuit that implements a logical diagonal Clifford operator up to Pauli correction (see [3, 4] and Supplementary Materials Lem. LABEL:supmat:lem:depth_1_diagonal_lifts). As there is always a Pauli operator with the appropriate (anti)commutation relations with the stabilizers and logical operators of the code to fix any logical/stabilizer action sign issues, we can ignore this subtlety [Sayginel2024].
The requirement that $\pi\tau$ is an involution guarantees that the upper-right block of (3) is symmetric, and thus corresponds to a valid diagonal Clifford operator. This generally restricts the number of automorphisms that may be leveraged to produce valid logical gates. Crucially, this is insignificant for the SHYPS codes we introduce in Section V, where we have sufficient symmetry to efficiently implement all diagonal operators in a code block.
V Code constructions and logical gates
The constructions (2) and (3) provide a framework for implementing logical Clifford operators with fault-tolerant, low-depth circuits. However the number of such operators that exist for a given subsystem CSS code $\mathcal{C}$ clearly scales with the size of ${\mathrm{Aut}}(\mathcal{C})$. This motivates a search for quantum codes with high degrees of permutation symmetry, to achieve the number of fixed-depth Clifford generators necessary for efficient compilation. There are many methods for constructing quantum codes [28, 29, 30, 31, 32, 33, 34], but in this work we focus on a subsystem hypergraph product construction that allows us to leverage known highly symmetric classical codes, to produce quantum code automorphisms: Let $H_{i}$ be parity check matrices for two classical $(n_{i},k_{i},d_{i})$-codes with $i=1,2$ and codespace $\ker H_{i}$. The subsystem hypergraph product code [15] (SHP), denoted $SHP(H_{1},H_{2})$, is the subsystem CSS code with gauge generators
$G_{X}=(H_{1}\otimes I_{n_{2}}),\ G_{Z}=(I_{n_{1}}\otimes H_{2}),$
and parameters $n_{1}n_{2},k_{1}k_{2},\min(d_{1},d_{2})] [$li2020numerical, 3.B].
Now the classical codes $\ker H_{i}$ have analogously defined automorphism groups, and crucially these lift to distinct automorphisms of $SHP(H_{1},H_{2})$.
Lemma V.1.
Let $(\sigma_{1},\sigma_{2})\in{\mathrm{Aut}}(\ker H_{1})\times{\mathrm{Aut}}(\ker H_{2})$. Then $\sigma_{1}\otimes\sigma_{2}\in{\mathrm{Aut}}(SHP(H_{1},H_{2}))$.
Note that a similar prescription for lifting classical symmetries has been recently proposed for the stabilizer hypergraph product codes (see [xu2024, App. A.1] and [hong2024, App. D]), in the restricted instance that the linear transformations of the classical checks $H_{i}$ induced by each $\sigma_{i}$, are themselves also permutations. By considering the subsystem code here, we remove such limitations and produce a far greater number of quantum code automorphisms (see Supplementary Materials Section VIII.6 for further details).
To capitalise on V.1, we pair the SHP construction with the highly symmetric classical simplex codes, referring to these as subsystem hypergraph product simplex (SHYPS) codes. A complete description of the SHYPS family (parameterized by integers $r\geq 3$) is given in Supplementary Materials VIII.5 but we note here that each instance, denoted $SHYPS(r)$ has parameters
$n(r),k(r),d(r)]=[(2^{r}-1)^{2},r^{2},2^{r-1}].$
Moreover, this is a QLDPC code family, as each $SHYPS(r)$ has weight-3 gauge generators. An immediate corollary of V.1 and [book:MW, Ch. 8.5] is that
$|{\mathrm{Aut}}(SHYPS(r))|\geq|GL_{r}(2)|^{2}=O(2^{2r^{2}}),$
which grows exponentially in the number of logical qubits $k(r)=r^{2}$, as required in Section III. By utilising these automorphisms with the constructions outlined in Section IV.2 we are able to efficiently generate all CNOT and diagonal Clifford operators in the SHYPS codes.
The logical action of operators (2) and (3) may be characterized explicitly: For any pair $g_{1},g_{2}\in GL_{r}(2)$ there exists a corresponding automorphism $\sigma_{1}\otimes\sigma_{2}\in{\mathrm{Aut}}(SHYPS(r))$ such that the logical cross-block CNOT operator
$\begin{bmatrix}I_{k}&g_{1}\otimes g_{2}&\\ 0&I_{k}&&\\ &&I_{k}&0\\ &&g_{1}^{-T}\otimes g_{2}^{-T}&I_{k}\\ \end{bmatrix}\in Sp_{4k}(2),$ (4)
is implemented by the depth-1 physical circuit of type (2). Furthermore, arbitrary logical CNOT circuits on $b$ blocks of $k$ logical qubits can be constructed from a sequence of $2bk(1+o(1))$ such operators (see Supplementary Materials Thm. LABEL:supmat:thm:cross-block-cnot-operators and Cor. LABEL:supmat:cor:b-block-cnot).
To characterize the logical action of diagonal operators (3) we first observe that the physical qubits of $SHYPS(r)$ may be naturally arranged in an $2^{r}-1\times 2^{r}-1$ array such that the reflection across the diagonal is a $ZX$-duality exchanging $G_{X}$ and $G_{Z}$ (see Supplementary Materials Lem. LABEL:supmat:lem:depth_1_diagonal_lifts). We similarly arrange the logical qubits in an $r\times r$ array, and denote the reflection that exchanges rows and columns by $\tau$. Then for all $g\in GL_{r}(2)$, there exists $\sigma\otimes\sigma^{T}\in{\mathrm{Aut}}(SHYPS(r))$ such that the logical diagonal Clifford operator
$\begin{bmatrix}I&\left(g\otimes g^{T}\right)\cdot\tau\\ 0&I\end{bmatrix}$ (5)
is implemented by a corresponding generator of type (3). The operators (5) have depth 1 and are fault-tolerant with circuit distance equal to the code distance. Moreover, they are alone sufficient to generate all logical diagonal Clifford operators on an SHYPS code block in depth at most $k(1+o(1))$ (Supplementary Materials Thm. LABEL:supmat:thm:DiagGensType1Generate).
For generation of the full logical Clifford group, we note that $SHYPS(r)$ possesses a Hadamard type [3] fold-transversal gate whereby the logical all-qubit Hadamard operator (up to logical SWAP) is implemented fault-tolerantly (see Supplementary Materials Lem. LABEL:supmat:lem:all_qubit_hadamard). By then applying a Clifford decomposition as discussed in Section III (see also Supplementary Materials LABEL:supmat:sec:SHSCompSummary), we bound the cost of implementing an arbitrary Clifford operator:
Theorem V.2.
Let $r\geq 3$. An arbitrary logical Clifford operator on $b$ blocks of the $SHYPS(r)$ code may be implemented fault-tolerantly in depth $4bk(r)(1+o(1))$.
In particular, the SHYPS codes achieve the desired $O(1)$ Clifford compiling ratio. Moreover this bound is competitive with best known depths of $2bk+O(log^{2}(bk))$ for compiling Cliffords on $bk$ unencoded qubits [17]. We achieve further reductions in overhead for logical permutations and arbitrary Hadamard circuits (see Supplementary Materials Tables LABEL:supmat:tab:cnot_depth, LABEL:supmat:tab:diagonal_depth and LABEL:supmat:tab:had_swap_depth for details).
VI Performance of the SHYPS Code
We present numerical simulations to evaluate the circuit-level noise performance of the SHYPS code family. Two types of simulations were performed. First, in section VI.1, memory simulations for two different instances of SHYPS codes are benchmarked against comparably scaled surface codes. Second, in section VI.2, we present a circuit-level noise logic simulation on two blocks of the $[49,9,4]$ SHYPS code. The logical Clifford circuit is randomly sampled from the $18$ qubit Clifford group and synthesized into low-depth generators using the techniques explained in Section V. For a more detailed treatment of the simulations presented in this section, see Supplementary Materials LABEL:supmat:fault-tolerant-demonstration.
VI.1 Memory Simulation
Figure 2: Simulation results for quantum memories under circuit-level noise for SHYPS and surface codes. For these simulations, only $Z$-type detectors are used.
Figure 2 shows the normalized logical error rate $p_{L}$ from memory simulations with $d$ syndrome extraction rounds. The logical error rates for the scaled surface codes are adjusted according to the number of logical qubits  [^7] Note that this is equivalent to considering multiple distance-$d$ surface code patches required to provide the same number of logical qubits as the distance $d$ SHYPS code. to demonstrate that the SHYPS code is competitive with the surface code while requiring fewer physical qubits per logical qubit. Specifically, we compare $[n,k,d]$ SHYPS codes with even $d$, to $k$ copies of $\underline{d}^{2},1,\underline{d}$]] rotated surface codes with uneven $\underline{d}=d-1$  [^8] The maximum weight of guaranteed correctable errors for distance $d$ and d1d-1 codes is the same when $d$ is even.. This comparison is shown in Table 2, where we have used the code pseudo-thresholds and logical error rates at $p=\{3\times 10^{-4},10^{-3}\}$ as the benchmarking metrics. The smallest member of the SHYPS family, the $[49,9,4]$ code, confers no pseudo-threshold advantages, but significantly outperforms the $[[81,9,3]]$ scaled distance-3 surface code in both physical qubit overhead (2x reduction) and logical error rates below pseudo-threshold. The other SHYPS code we simulated, the $[225,16,8]$ SHYPS code, is comparable in terms of performance to the $[[784,16,7]]$ scaled distance-7 surface code whilst reducing qubit count by $3.5\times$. It is true that the $[225,16,8]$ code has a lower pseudo threshold ($\approx 0.35\%$) than the $[[784,16,7]]$ scaled distance-7 surface code ($\approx 0.8\%$), but logical error rate decreases more aggressively in the sub-theshold regime: the slope of the logical error rates for both codes intersect at $p=0.05\%$, a point below which the $[[225,16,8]]$ SHYPS code outperforms the $[[784,16,7]]$ scaled distance-7 surface code. We anticipate that a decoder specifically tailored to SHYPS codes would further improve the pseudo-thresholds reported here.
Code $P_{L}(3\times 10^{-4})$ $P_{L}(10^{-3})$ $p_{th}$
$[49,9,4]$ $5\times 10^{-5}$ $1.2\times 10^{-3}$ $3.2\times 10^{-3}$
$[[81,9,3]]$ $2\times 10^{-4}$ $2.3\times 10^{-3}$ $3.8\times 10^{-3}$
$[225,16,8]$ $6.2\times 10^{-8}$ $7\times 10^{-4}$ $3.5\times 10^{-3}$
$[[784,16,7]]$ $3\times 10^{-7}$ $4\times 10^{-4}$ $8\times 10^{-3}$
Table 2: Logical error rates per round of error correction at physical error rates $3\times 10^{-4}\ (0.03\%)$, $10^{-3}\ (0.1\%)$ and pseudo-thresholds. Here, $P_{L}(p)$ denotes the block logical error rate per round of syndrome extraction at physical error rate pp. The pseudo-threshold is defined by the condition $P_{L}(p_{th}=p)=kp$, where $k$ is the number of logical qubits.
Most impressively, the SHYPS code simulations rely on a sliding window decoding approach with a small window size (2) and commit size (1). The use of a (2,1) sliding window decoder to achieve a high error correction performance is consistent with single-shot properties –the ability to decode based on a single syndrome extraction round between each logical operation. For a more detailed explanation about the single-shot properties of SHYPS codes, see Supplementary Materials LABEL:supmat:fault-tolerant-demonstration.
VI.2 Clifford Simulation
We now apply our decomposition and logical generator constructions in a simulation of a randomly sampled Clifford operator on 18 logical qubits in two code blocks of the $[49,9,4]$ SHYPS code, up to an in-block CNOT circuit. This prevents the need to use extra auxiliary code blocks to implement in-block CNOTs while ensuring all types of transversal operations appear in the simulation. Synthesizing the sampled Clifford using the $\mathrm{DZ}-\mathrm{CX}-\mathrm{DX}-\mathrm{DZ}(1)$ decomposition requires 63 fault-tolerant logical generators, and consequently a total of 126 logical generators to implement both the Clifford circuit and its inverse. The simulation proceeds as follows: initialize both code blocks in the encoded all-zero logical state; interleave each logical generator (both of the Clifford and its inverse) with gauge generator measurements; read out the state using a transversal $Z$ measurement.
Figure 3: Simulation results for a depth-$126$ logical quantum circuit and a memory composed from two code blocks of the $[49,9,4]$ SHYPS code. $Z$- and $X$-type detectors are used in both cases.
Figure 3 shows the normalized logical error rate $p_{L}$ for the described simulation. For comparison, we also include the logical error rate per syndrome extraction round for a $d$-round SHYPS quantum memory with $18$ logical qubits built from two blocks of the $[49,9,4]$ SHYPS code (adjusting SHYPS error rates to account for multiple codeblocks is analogous to scaling surface code data, see Supplementary Materials LABEL:supmat:fault-tolerant-demonstration for details). Simulation of the Clifford operator achieves near-memory error correction performance in terms of error suppression, demonstrating that logical operations can be executed efficiently and fault-tolerantly in SHYPS codes.
VII Conclusion
QLDPC codes promise to reduce physical qubit overheads compared to existing surface codes. Despite their advantage as memories, it has been unclear whether any multi-logical qubit code (QLDPC or otherwise) could execute logical operations with as much parallelism as a surface code. Were this to remain unresolved, it may have crippled the speed of any QLDPC-code-based quantum computer when running highly-parallel circuits. By using a product construction of highly-symmetric classical simplex codes alongside novel compiling methods, we have shown that the resultant SHYPS code family shatters this barrier, achieving the same asymptotic logical circuit depth as unencoded circuits under state-of-the-art algorithms. Remarkably, the resulting logical circuits retain strong fault-tolerance guarantees that are reflected in deep logical simulations showing near-memory performance even under circuit-level noise.
Two critical directions demand further exploration. Developing QEC codes with analogous properties but even better rates would enable further space improvements without making a time trade-off. More significantly, extending compiling parallelism to measurement parallelism would unlock for QLDPC codes every trick the surface code has at its disposal for reducing run-times via auxiliary code blocks. Without these advantages in parallelism, the reasons to consider using the surface code over QLDPC codes reduces to two factors: connectivity and simplicity. For any architecture where the necessary connectivity is achievable, it now seems all but certain: QLDPC codes are capable of driving down physical overheads without increasing time overheads, and, as a result of this work, appear to be the most compelling path to quantum computers that can perform commercially relevant algorithms.
Acknowledgements
We thank Polina Bychkova, Zach Schaller, Bogdan Reznychenko, and Kyle Wamer for their contributions to the development of simulation infrastructure.
Author contributions
A.J.M. and A.N.G. designed the quantum codes studied in this manuscript. The depth costing of logical operators was done by A.J.M., A.S. and A.N.G. P.F., D.C., J.R. and A.O.Q. designed the sliding window decoder and optimized decoder parameters used in the circuit-level numerical simulations. P.F., D.C., R.H., J.R., A.O.Q. and A.N.G. performed the numerical simulations and post-processed the resulting data. The software used for these simulations was designed and written by C.D., R.H., A.E., P.F., J.R., and D.C.. A.J.M., A.N.G., P.F., D.C., A.S., J.R., A.O.Q., R.H., S.J.B., N.R.L.-H. and S.S. contributed to writing and editing the manuscript.
Competing interests
US Patent Application 19/263,208 (filed 8 July 2025, naming A.J.M and A.N.G. as co-inventors), PCT Application PCT/IB2025/057033 (filed 11 July 2025, naming A.J.M and A.N.G. as co-inventors), US Provisional Patent Application 63/720,973 (filed on November 15, 2024, naming A.N.G. and A.S as co-inventors), and US Provisional Patent Application 63/764,648 (filed 28 February 2025, naming A.N.G as the inventor) contain technical aspects from this paper.
Additional information
Supplementary Information is available for this paper. Correspondence and requests for materials should be addressed to Stephanie Simmons at ssimmons@photonic.com.
Data availability
The simulation software to generate data reported in this paper is available at:
https://github.com/PhotonicInc/ComputingEfficientlyInQLDPCCodes.
Supplementary Material
The supplementary material is broken into 3 parts. First, in Section VIII we survey the necessary background information to introduce our new code family, the subsystem hypergraph product simplex codes (SHYPS) codes. In particular, we discuss the notion of code automorphisms, and how the subsystem hypergraph product construction yields highly symmetric quantum codes, from well chosen classical inputs.
Next, in Sections IX-LABEL:supmat:sec:_Hadamard we demonstrate how automorphisms of the SHYPS codes can be leveraged to obtain low-depth implementations of logical Clifford operators. Taking each of the CNOT, diagonal, and Hadamard-SWAP families in turn, we produce fault-tolerant generators (typically physical depth 1), and demonstrate efficient compilation of arbitrary operators. Utilising a novel decomposition of Clifford operators, these results are combined in LABEL:supmat:sec:SHSCompSummary to yield an overarching bound on the depth of Clifford implementations in SHYPS codes. We refer the reader to Tables LABEL:supmat:tab:cnot_depth, LABEL:supmat:tab:diagonal_depth and LABEL:supmat:tab:had_swap_depth, for a detailed summary of results.
Lastly, in LABEL:supmat:fault-tolerant-demonstration we discuss syndrome extraction and decoding of the SHYPS codes, and we provide details about our numerical simulations.
VIII Mathematical preliminaries and code constructions
We begin the preliminaries section with a review of the Pauli and Clifford groups, including the binary symplectic representation and examples of key types of Clifford operations that will be the focus of later sections.
VIII.1 Review of Paulis and Cliffords
The Pauli group on $n$ qubits is defined as
$\mathcal{P}_{n}$ $:=$ $\langle\mathrm{i},X_{j},Z_{j}\mid j\in\{1,\dots,n\}\rangle$
$=$ $\langle X_{j},Y_{j},Z_{j}\mid j\in\{1,\dots,n\}\rangle.$
For many applications, it is convenient to ignore global phases involved in Pauli operators and instead consider elements of the phaseless Pauli group, $\mathcal{P}_{n}/\left<\mathrm{i}\right$>. The phaseless Pauli group is abelian and has order $4^{n}$. Additionally, every non-trivial element has order 2, and hence $\mathcal{P}_{n}/\left<\mathrm{i}\right>\cong\mathbb{F}_{2}^{2n}$. This isomorphism can be made explicit in the following manner: as $XZ=\mathrm{i}Y$, any $P\in\mathcal{P}_{n}/\left<\mathrm{i}\right$> may be written uniquely as
$\prod_{i=1}^{n}X_{i}^{u_{i}}Z_{i}^{v_{i}}=:X^{u}Z^{v},$
where $u=(u_{1},\dots,u_{n})\in\mathbb{F}_{2}^{n}$, and $v$ is defined analogously. The combined vector $(u\mid v)\in\mathbb{F}_{2}^{2n}$ is known as the binary symplectic representation of $P$. We equip $\mathbb{F}_{2}^{2n}$ with a symplectic form $\omega:\mathbb{F}_{2}^{2n}\times\mathbb{F}_{2}^{2n}\longrightarrow\mathbb{F}_{2}$ such that for $(u\mid v),(s\mid t)\in\mathbb{F}_{2}^{2n}$,
$(u\mid v),(s\mid t)$ $\mapsto$ $ut^{T}+sv^{T}.$
This form captures the (anti-)commutativity of Paulis (that is lost when global phases are ignored), as $X^{u}Z^{v}$ and $X^{s}Z^{t}$ anti-commute if and only if
$\omega((u\mid v),(s\mid t))=1.$
For example in $\mathcal{P}_{2}$, the element $X_{1}$ anti-commutes with $Y_{1}Z_{2}$, and we check that
$X_{1}$ $\mapsto$ $(1,0,0,0)$
$Y_{1}Z_{2}\sim X_{1}Z_{1}\cdot Z_{2}$ $\mapsto$ $(1,0,1,1),$
and
$\omega((1,0,0,0),(1,0,1,1))=1.$
The Clifford group on $n$ qubits, denoted $\mathrm{Clif}_{n}$, is the normalizer of the $n$-qubit Pauli group within $U(n)$. I.e.,
$\mathrm{Clif}_{n}$ $:=$ $N_{U(n)}(\mathcal{P}_{n})$
$=$ $\{g\in U(n)\,:\,gPg^{-1}\in\mathcal{P}_{n}\text{ for all }P\in\mathcal{P}_{n}\}\,.$
For example, the two-qubit controlled-not operator $CNOT_{i,j}$, the single-qubit phase gate $S_{i}$, and the single-qubit Hadamard gate $H_{i}$, are all Clifford operators. Further examples include all gates of the form $e^{\mathrm{i}\theta}I$, but as the global phase is typically unimportant we restrict our attention to
$\mathcal{C}_{n}:=\mathrm{Clif}_{n}/\{e^{\mathrm{i}\theta}\}\times\left<\frac{1+\mathrm{i}}{\sqrt{2}}\right$>.
We likewise refer to this as the Clifford group, and note that
$\mathcal{C}_{n}=\left<CNOT_{i,j},S_{i},H_{i}\,:\,1\leq i,j\leq n\right$>.
The action of a Clifford operator via conjugation corresponds to a linear transformation of $\mathbb{F}_{2}^{2n}$ in the binary symplectic representation. Moreover, as conjugation preserves the (anti-)commutation of of Pauli operators, the corresponding linear transformations preserve the symplectic form. The collection of such linear transformations is known as the symplectic group and is denoted $Sp_{2n}(2)$. Finally, observe that $\mathcal{P}_{n}\leq\mathcal{C}_{n}$ but that conjugation by a Pauli operator induces at most a change of sign, which is ignored by the binary symplectic representation. Taking the quotient by this trivial action, we see that $\mathcal{C}_{n}/\mathcal{P}_{n}\cong Sp_{2n}(2)$ [Rengaswamy2018, Thm. 15].
The representation of Clifford operators by $2n\times 2n$ binary matrices is key to the efficient simulation of stabilizer circuits [aaronson2004]. Moreover, as we’ll demonstrate, it is a useful framework for synthesising efficient implementations of logical gates in quantum codes (see also [Rengaswamy2018]).
We conclude this section with an explicit construction of some symplectic representations for Clifford operators. Following [aaronson2004], we assume that matrices in $Sp_{2n}(2)$ act on row vectors from the right. I.e., given $P\in\mathcal{P}_{n}$ with binary symplectic representation $(u\mid v)$, and $g\in\mathcal{C}_{n}/\mathcal{P}_{n}$ with binary symplectic representation $G$, we have $gPg^{-1}\longleftrightarrow(u\mid v)G$. In particular, the images of the Pauli basis $X_{i},Z_{i}$, are given by the $i$th and $(i+n)$th rows of $G$, respectively.
Example VIII.1.
(CNOT circuits) As CNOT circuits map $X$-type Paulis to other $X$-type Paulis (and similar for $Z$-type), in $Sp_{2n}(2)$ they have the form
$\{\begin{pmatrix}C&0\\ 0&C^{-T}\end{pmatrix}\,:\,C\in GL_{n}(2)\}.$
For simplicity we typically describe such operators solely by the matrix $C$.
For example, $CNOT_{1,2}\in\mathcal{C}_{3}$ has $C$ defined as follows
$\begin{pmatrix}1&1&0\\ 0&1&0\\ 0&0&1\end{pmatrix}\in GL_{3}(2).$
Example VIII.2.
(Diagonal Clifford operators) The Clifford operators that act diagonally on the computational basis form an abelian group, generated by single-qubit phase gates $S_{i}$ and the two-qubit controlled-Z gate $CZ_{i,j}$. Hence, modulo Paulis, they are represented by symplectic matrices of the form
$\{\begin{pmatrix}I&B\\ 0&I\end{pmatrix}\,:\,B\in M_{k}(2),B^{T}=B\},$
where the diagonal and off-diagonal entries of the symmetric matrix $B$ determine the presence of $S$ and CZ gates, respectively.
For example, $S_{1}\cdot CZ_{1,2}\in\mathcal{C}_{2}$ corresponds to
$\begin{pmatrix}1&0&1&1\\ 0&1&1&0\\ 0&0&1&0\\ 0&0&0&1\end{pmatrix}\in Sp_{4}(2).$
And the action of this example on a Pauli, $X_{1}$,
$(S_{1}\cdot CZ_{1,2})\cdot X_{1}\cdot(S_{1}\cdot CZ_{1,2})^{-1}=Y_{1}Z_{2}$
corresponds to
$(1,0,0,0)\cdot\begin{pmatrix}1&0&1&1\\ 0&1&1&0\\ 0&0&1&0\\ 0&0&0&1\end{pmatrix}=(1,0,1,1).$
Example VIII.3.
(Hadamard circuits) In $Sp_{2n}(2)$, Hadamard circuits $H^{v}=\prod H_{i}^{v_{i}}$ have the form
$\{\begin{pmatrix}I_{n}+\text{diag}(v)&\text{diag}(v)\\ \text{diag}(v)&I_{n}+\text{diag}(v)\\ \end{pmatrix}\,:\,v\in\mathbb{F}_{2}^{n}\},$
where $\text{diag}(v)$ is the diagonal matrix with entries $v_{1},\dots,v_{n}$.
For example, the transversal Hadamard operator $H^{\otimes 3}=H^{(1,1,1)}\in\mathcal{C}_{3}$ corresponds to
$\begin{pmatrix}0&I_{3}\\ I_{3}&0\end{pmatrix}\in Sp_{6}(2).$
VIII.2 Subsystem codes
A quantum stabilizer code on $n$ physical qubits is the common +1+1 eigenspace of a chosen abelian subgroup $\mathcal{S}\leq\mathcal{P}_{n}$ with $I\not\in\mathcal{S}$. The subgroup $\mathcal{S}$ is known as the stabilizer group, and moreover if $\mathcal{S}$ admits a set of generators that are either $X$-type or $Z$-type Pauli strings, then the code is called CSS [Calderbank1996]. Quantum subsystem codes are the natural generalisation of stabilizer codes, in that they are defined with respect to a generic subgroup $\mathcal{G}\leq\mathcal{P}_{n}$ known as the gauge group [Poulin2005]. Moreover, subsystem codes are typically interpreted as the subsystem of a larger stabilizer code whereby a subset of the logical qubits are chosen to not store information and the action of the corresponding logical operators is ignored. More formally, given gauge group $\mathcal{G}$, the corresponding stabilizer group $\mathcal{S}$ is the centre of $\mathcal{G}$ modulo phases
$\left<\mathcal{S},iI\right>=Z(\mathcal{G}):=C_{\mathcal{P}_{n}}(\mathcal{G})\cap\mathcal{G}.$
Phases are purposefully excluded (in particular, $I\notin\mathcal{S}$) to ensure the fixed point space of $\mathcal{S}$ is nontrivial, and said space decomposes into a tensor product $\mathcal{C}_{\mathcal{L}}\otimes\mathcal{C}_{\mathcal{G}}$, where elements of $\mathcal{G}\backslash\mathcal{S}$ fix only $\mathcal{C}_{\mathcal{L}}$. The subspaces $\mathcal{C}_{\mathcal{L}}$ and $\mathcal{C}_{\mathcal{G}}$ are said to contain the logical qubits, and gauge qubits, respectively.
The logical operators of the subsystem code are differentiated into two types, based on their action on $\mathcal{C}_{\mathcal{L}}\otimes\mathcal{C}_{\mathcal{G}}$: those given by $C_{\mathcal{P}_{n}}(\mathcal{G})\backslash\mathcal{G}$ that act nontrivially only on $\mathcal{C}_{\mathcal{L}}$ are known as bare logical operators. Whereas operators acting nontrivially on both $\mathcal{C}_{\mathcal{L}}$ and $\mathcal{C}_{\mathcal{G}}$ are known as dressed logical operators, and are given by $C_{\mathcal{P}_{n}}(\mathcal{S})\backslash\mathcal{G}$. Note that a dressed logical operator is a bare logical operator multiplied by an element of $\mathcal{G}\backslash\mathcal{S}$. The minimum distance $d$ of the subsystem code is the minimal weight Pauli that acts nontrivially on the logical qubits $\mathcal{C}_{\cal{L}}$, i.e., the minimum weight of a dressed logical operator
$d=\min\{|P|\,:\,P\in C_{\mathcal{P}_{n}}(\mathcal{S})\backslash\mathcal{G}\}.$
We say that a subsystem code is an $[n,k,d]$ code if it uses $n$ physical qubits to encode $k$ logical qubits with distance $d$. The notation $[n,k,g,d]$ that additionally indicates the number of gauge qubits $g$, is also commonplace (but unused throughout this paper).
The advantages of subsystem codes are most evident when the gauge group $\mathcal{G}$ has a generating set composed of low-weight operators but the stabilizer group $\mathcal{S}$ consists of high-weight Paulis. Measuring the former operators requires circuits of lower depth (therefore reducing computational overheads), and the measurement results can be aggregated to infer the stabilizer eigenvalues. In fact, for the codes considered here, the difference in the stabilizer weights and gauge operator weights is a factor of the code distance (see LABEL:supmat:fault-tolerant-demonstration for more details). We note that it is exactly the measurement of operators in $\mathcal{G}\backslash\mathcal{S}$ that act nontrivially on $\mathcal{C}_{\mathcal{G}}$, that prevents the gauge qubits from storing information during computation, as these measurements impact the state of $\mathcal{C}_{\mathcal{G}}$.
In this work we restrict our attention to $[n,k,d]$-subsystem codes that are also of CSS type, with $X$- and $Z$-type gauge generators determined by matrices $G_{X}\in\mathbb{F}_{2}^{r_{X}\times n}$ and $G_{Z}\in\mathbb{F}_{2}^{r_{Z}\times n}$, respectively. Here each vector $v\in\mathbb{F}_{2}^{n}\cap\text{RowSpace}(G_{X})$ denotes an $X$-type gauge operator $X^{v}$, with $Z$-type gauge generators similarly defined. The associated stabilizers will also be of CSS type, and denoted by $S_{X},S_{Z}$.
VIII.3 Automorphisms of codes
Code automorphisms are a promising foundation for computing in QLDPC codes as they can provide nontrivial logical gates implementable by permuting, or in practice simply relabelling, physical qubits. In this section, we review permutation automorphisms of classical and quantum codes, which will serve as the backbone of our logical operation constructions.
Let’s first set some notation: given a permutation $\sigma\in S_{n}$ of the symbols $\{1,2,\dots,n\}$ in cycle notation, we identify $\sigma$ with the permutation matrix whose $(i,j)$th entry is 1 if $i=\sigma(j)$, and zero otherwise. For example
$(1,2,3)(4,5)\in S_{5}\mapsto\begin{pmatrix}0&0&1&0&0\\ 1&0&0&0&0\\ 0&1&0&0&0\\ 0&0&0&0&1\\ 0&0&0&1&0\end{pmatrix}$.
The permutation defined above maps the initial indices $\{1,2,3,4,5\}$ to $\{2,3,1,5,4\}$. So given a standard basis vector $e_{i}\in\mathbb{F}_{2}^{n}$, $\sigma$ acts on row vectors on the right as $e_{i}\sigma=e_{\sigma^{-1}(i)}$, and on column vectors on the left as $\sigma e_{i}^{T}=e_{\sigma(i)}^{T}$.
Definition VIII.4.
Let $C$ be an $(n,k,d)$-classical linear code described by a generator matrix $G\in\mathbb{F}_{2}^{k\times n}$. Then the (permutation) automorphism group ${\mathrm{Aut}}(C)$ is the collection of permutations $\sigma\in S_{n}$ that preserve the codespace. I.e., $\sigma\in{\mathrm{Aut}}(C)$ if for all $c=(c_{1},\dots,c_{n})\in C$,
$c\sigma=(c_{\sigma(1)},\dots,c_{\sigma(n)})\in C.$ (6)
As $C$ is linear, it suffices to check (6) on the basis given by the rows of $G$. Hence $\sigma\in{\mathrm{Aut}}(C)$ if and only if there exists a corresponding $g_{\sigma}\in GL_{k}(2)$ such that $G\sigma=g_{\sigma}G$ [book:MW, Lem. 8.12]. In particular, $g_{\sigma}$ represents the invertible linear transformation of the $k$ logical bits, induced by the permutation.
The parity checks of the code $C$ are similarly transformed by automorphisms: letting $C^{\perp}$ denote the dual classical code, we have ${\mathrm{Aut}}(C)={\mathrm{Aut}}(C^{\perp})$ [book:MW, Sec. 8.5]. So given a parity check matrix $H\in\mathbb{F}_{2}^{(n-k)\times n}$ for $C$ we have equivalently that $\sigma\in{\mathrm{Aut}}(C)$ if there exists $h_{\sigma}\in GL_{n-k}(2)$ such that $H\sigma=h_{\sigma}H$.
This definition extends naturally to quantum codes:
Definition VIII.5.
Let $C$ be an $[n,k,d]$ (CSS) subsystem code with $X$ and $Z$ type gauge generators determined by $G_{X}\in\mathbb{F}_{2}^{r_{X}\times n}$ and $G_{Z}\in\mathbb{F}_{2}^{r_{Z}\times n}$, respectively. Then ${\mathrm{Aut}}(C)$ consists of permutations that preserve the gauge generators, i.e., $\sigma\in S_{n}$ such that
$g_{\sigma,X}G_{X}=G_{X}\sigma,\ \,g_{\sigma,Z}G_{Z}=G_{Z}\sigma,$
for some $g_{\sigma,X}\in GL_{r_{X}}(2)$ and $g_{\sigma,Z}\in GL_{r_{Z}}(2)$.
The logical gate implemented by $\sigma\in{\mathrm{Aut}}(C)$ is determined by the permutation action on the code’s logical Pauli operators. In particular, $\sigma$ always gives rise to a permutation of the logical computational basis states of $C$ corresponding to a logical CNOT circuit [Grassl2013, Thm. 2].
In later sections we outline quantum constructions utilising classical codes, and consequently how classical automorphisms may be leveraged to produce quantum logical gates. In these instances, the logical CNOT implemented by the quantum code automorphism is a function of the associated classical linear transformations. This motivates an investigation of classical codes with high degrees of symmetry.
VIII.4 Classical simplex codes
Let $r\geq 3$ and define $n_{r}=2^{r}-1$, $d_{r}=2^{r-1}$. The classical simplex codes, denoted $C(r)$ are a family of $(n_{r},r,d_{r})$-linear codes, that are dual to the well-known Hamming codes. More specifically, we consider $C(r)$ with respect to a particular choice of parity check matrix: for each $3\leq r<500$  [^9] The upper bound of $500$ clearly encapsulates all codes that will ever be used in practice. That the result in fact holds for all $r$ is an open conjecture [blake96], there exists a three term polynomial $h(x)=1+x^{a}+x^{b}\in\mathbb{F}_{2}[x]/\left<x^{n_{r}}-1\right$> such that $\gcd(h(x),x^{n_{r}}-1)$ is a primitive polynomial of degree $r$ [blake96]. Then the $n_{r}\times n_{r}$ matrix
$H=\begin{pmatrix}h(x)\\ xh(x)\\ \vdots\\ x^{n_{r}-1}h(x)\end{pmatrix}$
is a parity check matrix (PCM) for $C(r)$ [book:MW, Lem. 7.5], where here we adopt the usual polynomial notation for cyclic matrices
$\sum_{i=0}^{n_{r}-1}a_{i}x^{i}\mod x^{n_{r}}-1\mapsto(a_{0},a_{1},\dots,a_{n_{r}-1})\in\mathbb{F}_{2}^{n_{r}}.$
Note there are many alternative choices for the PCM of $C(r)$; in fact $H$ chosen here has greater than $n_{r}-r$ rows and so this description contains redundancy. However what the choice above guarantees is that each row and column of $H$ has weight 3, leading to low-weight gauge generators and optimal syndrome extraction scheduling, of an associated quantum code (see LABEL:supmat:fault-tolerant-demonstration).
The simplex codes are examples of highly symmetric classical codes with large automorphism groups.
Lemma VIII.6.
The automorphism group of the simplex codes are as follows:
${\mathrm{Aut}}(C(r))\cong GL_{r}(2).$
Intuitively, VIII.6 means that each invertible linear transformation $g\in GL_{r}(2)$ of the $r$ logical bits, is implemented by a distinct permutation $\sigma$ of the $n_{r}$ physical bits [book:MW, Ch. 8.5].
Example VIII.7.
Let $r=3$. The polynomial $h(x)=1+x^{2}+x^{3}$ is primitive and hence the overcomplete parity check matrix
$H=\begin{pmatrix}1&0&1&1&0&0&0\\ 0&1&0&1&1&0&0\\ 0&0&1&0&1&1&0\\ 0&0&0&1&0&1&1\\ 1&0&0&0&1&0&1\\ 1&1&0&0&0&1&0\\ 0&1&1&0&0&0&1\end{pmatrix}$
defines the $(7,3,4)$-simplex code. A basis for $C(3)$ is given by the rows of generator matrix
$G=\begin{pmatrix}1&0&1&1&1&0&0\\ 0&1&0&1&1&1&0\\ 0&0&1&0&1&1&1\end{pmatrix}$, (7)
and we observe that
$\begin{pmatrix}1&1&0\\ 0&1&0\\ 0&0&1\end{pmatrix}\cdot G=G\cdot(2,4)(5,6).$
I.e., the bit permutation $(2,4)(5,6)\in{\mathrm{Aut}}(C(3))$ induces the linear transformation
$(v_{1},v_{2},v_{3})\mapsto(v_{1}+v_{2},v_{2},v_{3})$
on the 3 logical bits $(v_{1},v_{2},v_{3})\in\mathbb{F}_{2}^{3}$.
In the remainder of this work, we assume that all parity check matrices $H\in\mathbb{F}_{2}^{n_{r}\times n_{r}}$ for the classical simplex code are taken as above.
VIII.5 Subsystem hypergraph product simplex (SHYPS) codes
Here we describe our main quantum code construction, namely the subsystem hypergraph product simplex code, in greater detail.
Definition VIII.8.
Let $r\geq 3,n_{r}=2^{r}-1,d_{r}=2^{r-1}$, and let $H_{r}$ be the parity check matrix for the $(n_{r},r,d_{r})$-classical simplex code. Then the subsystem hypergraph product of two copies of $H_{r}$, denoted $SHYPS(r)$, is the subsystem CSS code with gauge generators
$G_{X}=(H\otimes I_{n_{r}}),\ G_{Z}=(I_{n_{r}}\otimes H).$ (8)
We call $SHYPS(r)$ the subsystem hypergraph product simplex code.
It’s clear by definition that the row/column weights of $G_{X}$ and $G_{Z}$ match those of $H$, and so the $SHYPS(r)$ codes are a QLDPC code family with gauge generators of weight 3. Moreover, the gauge generators have a particular geometric structure: we arrange the $n_{r}^{2}$ physical qubits of $SHYPS(r)$ in an $n_{r}\times n_{r}$ array with row major ordering, such that for standard basis vectors $\left<e_{1},\dots,e_{n_{r}}\right>=\mathbb{F}_{2}^{n_{r}}$, the vector $e_{i}\otimes e_{j}$ corresponds to the $(i,j)$th position of the $n_{r}\times n_{r}$ qubit array. For example, the first row of $H$ given in Example VIII.7 is $r_{1}=e_{1}+e_{3}+e_{4}$ and hence the first row $r_{1}\otimes e_{1}\in G_{X}$ indicates an $X$-type gauge generator supported on the vector
$r_{1}\otimes e_{1}=e_{1}\otimes e_{1}+e_{3}\otimes e_{1}+e_{4}\otimes e_{1},$
i.e., on the $(1,1),(3,1)$ and $(4,1)$ qubits of the array. It follows that gauge generators in $G_{X}$ (respectively $G_{Z}$) are supported on single columns (respectively rows) of the qubit array.
We follow [li2020numerical, Sec. 3B] to describe the parameters $[n,k,d]$ of $SHYPS(r)$: firstly recall from the above that $n=n_{r}^{2}$. Next, to calculate the number of encoded qubits $k$, observe that the Pauli operators that commute with all gauge generators, are generated by
$\mathcal{L}_{X}=(I_{n_{r}}\otimes G),\ \mathcal{L}_{Z}=(G\otimes I_{n_{r}}).$ (9)
Here $G$ is a chosen generator matrix for the classical simplex code (so in particular $G$ is a matrix of rank $r$ such that $HG^{T}=0$).
The centre of the gauge group $\left<G_{X},G_{Z}\right$> determines the stabilizers of a subsystem code. In particular, for $SHYPS(r)$ these are generated by
$S_{X}=(H\otimes G),\ S_{Z}=(G\otimes H).$ (10)
Finally, $k$ is calculated by comparing the ranks of $\mathcal{L}_{X}$ and $S_{X}$:
[…]
$=$ $n_{r}\cdot r-(n_{r}-r)\cdot r$
$=$ $r^{2}.$
More specifically, $\text{RowSpan}(\mathcal{L}_{X}\backslash S_{X})$ determines the space of logical $X$-operators, with logical $Z$-operators defined similarly. However, as indicated in Section VIII.2, we consider the action of these operators up to multiplication by the gauge group. Hence the minimum distance $d$ of the code is given by the minimum weight operator in $\left<\mathcal{L}_{X},\mathcal{L}_{Z}\right>\cdot\left<G_{X},G_{Z}\right>-\left<G_{X},G_{Z}\right$>. For subsystem hypergraph product codes, this is exactly the minimum distance of the involved classical codes, and hence $d(SHYPS(r))=d_{r}=2^{r-1}$ [li2020numerical, Sec. 3B].
In summary, we have the following
Theorem VIII.9.
Let $r\geq 3$ and let $H$ be the parity check matrix for the $(2^{r}-1,r,2^{r-1})$-classical simplex code described in Section VIII.4. The subsystem hypergraph product simplex code $SHYPS(r)$ is an $[n,k,d]$-quantum subsystem code with gauge group generated by $3$-qubit operators and
$n$ $=$ $(2^{r}-1)^{2},$
$k$ $=$ $r^{2},$
$d$ $=$ $2^{r-1}.$
It’s evident from the tensor product structure of (9) that like the physical qubits, the logical qubits of $SHYPS(r)$ may be arranged in an $r\times r$ array, such that logical operators have support on lines of qubits. In fact, recent work [Quintavalle2023] demonstrates that a basis of logical operators may be chosen such that pairs of logical $X/Z$ operators have supports intersecting in at most one qubit. The following result is an immediate application of [Quintavalle2023, Thm. 1]:
Theorem VIII.10.
Let $r\geq 3$. There exists a generator matrix $G\in\mathbb{F}_{2}^{r\times n_{r}}$ for the classical simplex code and an ordered list of $r$ bit indices $\pi(G)\subset[1..n_{r}]$ known as pivots, such that $PG^{T}=I_{r}$ for the pivot matrix $P\in\mathbb{F}_{2}^{r\times n_{r}}$: $P_{i,j}=1$ if and only if $\pi(G)[i]=j$. Moreover, the matrices
$L_{X}=(P\otimes G),\ L_{Z}=(G\otimes P),$
form a symplectic basis for the logical $X/Z$ operators of $SHYPS(r)$.
The proof of [Quintavalle2023, Thm. 1] is constructive and, as the name pivots suggests, relies on a modified version of the Gaussian elimination algorithm [Quintavalle2023, Alg. 1]. This yields a so-called strongly lower triangular basis for the simplex code, represented by $G$, with rows $\{g_{i}\}_{i\in\pi(G)}$ indexed by the pivots. As $P$ has row weights equal to one, we see that the matrix products $PG^{T}=I_{r}$ and $L_{X}L_{Z}^{T}=I_{r^{2}}$ hold not only over $\mathbb{F}_{2}$ but over $\mathbb{R}$. Hence, given pairs of pivots $(i,j),(k,l)\in\pi(G)^{2}$, the associated logical Paulis $\overline{X}_{i,j},\overline{Z}_{k,l}$ given by basis vectors $e_{i}\otimes g_{j}\in L_{X}$ and $g_{k}\otimes e_{l}\in L_{Z}$ have intersecting support if and only if $(i,j)=(l,k)$. Moreover, this intersection is on the $(i,j)$th qubit of the array. From this point, we assume that $G$, and the logical operators of $SHYPS(r)$, are of the form above.
VIII.6 Lifting classical automorphisms
The geometric structure of $SHYPS(r)$ suggests a natural way to lift automorphisms of the simplex code by independently permuting either rows or columns of the qubit array. In this manner, we see that $SHYPS(r)$ inherits $|GL_{r}(2)|^{2}=O(2^{2r^{2}})$ automorphisms from the two copies of the classical simplex code, and hence ${\mathrm{Aut}}(SHYPS(r))$ grows exponentially with the number of logical qubits $k=r^{2}$.
Lemma VIII.11.
Let $r\geq 3$ and $C$ be the $(n_{r},r,d_{r})$-simplex code with automorphisms $\sigma_{1},\sigma_{2}\in{\mathrm{Aut}}(C)$. Then $\sigma_{1}\otimes\sigma_{2}\in{\mathrm{Aut}}(SHYPS(r))$. Furthermore,
$|{\mathrm{Aut}}(SHYPS(r))|\geq|GL_{r}(2)|^{2}.$
Proof.
Clearly $\sigma_{1}\otimes\sigma_{2}$ is a permutation of the required number of qubits $n=n_{r}^{2}$ and so we need only check that it preserves the gauge generators $G_{X}$ and $G_{Z}$. By definition of code automorphisms, for each $\sigma_{i}\in{\mathrm{Aut}}(C)$ there exists a corresponding $h_{\sigma_{i}}\in GL_{n_{r}}(2)$ such that $h_{\sigma_{i}}H=H\sigma_{i}$. Hence
$G_{X}\cdot(\sigma_{1}\otimes\sigma_{2})$ $=$ $(H\sigma_{1}\otimes\sigma_{2})$
$=$ $(h_{\sigma_{1}}H\otimes\sigma_{2})$
$=$ $(h_{\sigma_{1}}\otimes\sigma_{2})\cdot G_{X},$
with $G_{Z}$ similarly preserved. For the second claim, note that a tensor product of matrices $A\otimes B$ is the identity if and only if $A$ and $B$ are also identity. Hence each pair $\sigma_{1},\sigma_{2}\in{\mathrm{Aut}}(C_{1})\times{\mathrm{Aut}}(C_{2})$ produces a distinct $\sigma_{1}\otimes\sigma_{2}\in{\mathrm{Aut}}(SHYPS(r))$. ∎
Note that the above result naturally generalises to all subsystem hypergraph product codes, constructed from possibly distinct classical codes, but it does not generalise to the stabilizer hypergraph product construction (HGP). A similar prescription for lifting classical automorphisms is available for HGP codes [xu2024, App. A.1] and [hong2024, App. D], but only in the restricted instance that each $h_{\sigma_{i}}$ is also a permutation (see [xu2024, Diagram A.3], where $h_{\sigma_{i}}$ are denoted $\gamma_{0}^{i}$). By considering the subsystem hypergraph product in this work, we are able to lift all $|GL_{r}(2)|$ classical simplex code automorphisms.
To determine the logical gate induced by the above automorphisms we examine the permutation action on the logical basis given by VIII.10.
Lemma VIII.12.
Let $r\geq 3$ and $C(r)$ be the $(n_{r},r,d_{r})$-simplex code with generator matrix $G$. Furthermore, assume that $\sigma_{1},\sigma_{2}\in{\mathrm{Aut}}(C)$ with corresponding linear transformations $g_{\sigma_{1}},g_{\sigma_{2}}\in GL_{r}(2)$ such that $g_{\sigma_{i}}G=G\sigma_{i}$. Then $\sigma_{1}\otimes\sigma_{2}\in{\mathrm{Aut}}(SHYPS(r))$ induces the following action on the basis of logical operators
$L_{Z}\cdot(\sigma_{1}\otimes\sigma_{2})=(g_{\sigma_{1}}\otimes g_{\sigma_{2}}^{-T})\cdot L_{Z},$
$L_{X}\cdot(\sigma_{1}\otimes\sigma_{2})=(g_{\sigma_{1}}^{-T}\otimes g_{\sigma_{2}})\cdot L_{X}.$
Proof.
For ease of presentation, let’s consider the action of $\sigma\otimes I$ where $\sigma\in{\mathrm{Aut}}(C)$ (the general case follows identically). Firstly observe that
$L_{Z}(\sigma\otimes I)=(G\sigma\otimes P)=(g_{\sigma}G\otimes P)=(g_{\sigma}\otimes I)L_{Z}.$
The permutation action on the basis of $X$-logicals is then given by
$L_{X}(\sigma\otimes I)=(P\sigma\otimes G),$
where
$I_{r^{2}}=L_{X}\cdot L_{Z}^{T}$ $=$ $L_{X}(\sigma\otimes I)\cdot(\sigma^{T}\otimes I)L_{Z}^{T}$
$=$ $(P\sigma\otimes G)\cdot(L_{Z}(\sigma\otimes I))^{T}$
$=$ $(P\sigma\otimes G)\cdot(G^{T}\otimes P^{T})\cdot(g_{\sigma}^{T}\otimes I).$
Collecting terms and noting that $A\otimes B=I_{r^{2}}$ if and only if both components are identity (recall also that $PG^{T}=I_{r}$), it follows that $P\sigma G^{T}=g_{\sigma}^{-T}.$ Now $H$ spans $\ker G$ and hence all solutions to the above are of the form
$P\sigma=g_{\sigma}^{-T}P+AH,$
for some $A\in M_{n_{r}}$. In summary, there exists $A$ such that
$L_{X}(\sigma\otimes I)$ $=$ $P\sigma\otimes G$
$=$ $(g_{\sigma}^{-T}P+AH)\otimes G$
$=$ $(g_{\sigma}^{-T}\otimes I)(P\otimes G)+(A\otimes I_{r})(H\otimes G)$
$=$ $(g_{\sigma}^{-T}\otimes I)L_{X}+(A\otimes I_{r})S_{X}.$
That is, up to stabilizers (the exact stabilizer determined by $A$), $\sigma\otimes I$ has the desired action on $L_{X}$. ∎
As a logical Clifford operator is determined (up to a phase) by its action on the basis of logical Paulis, Lemma VIII.12 demonstrates that the qubit permutation $\sigma_{1}\otimes\sigma_{2}$ induces the logical CNOT circuit corresponding to $g_{\sigma_{1}}^{-T}\otimes g_{\sigma_{2}}\in GL_{r^{2}}(2)$.
IX CNOT operators in SHYPS codes
In this section, the collection of automorphisms
$\{\sigma_{1}\otimes\sigma_{2}\mid\sigma_{i}\in{\mathrm{Aut}}(C(r))\}\subset{\mathrm{Aut}}(SHYPS(r))$
serves as the foundation for generating all logical CNOT operators in the SHYPS codes.
First, following [Grassl2013], we show how all logical cross-block CNOT operators (with all controls in a first code block, and all targets in a second code block) may be attained by sequences of physical depth-1 circuits, that interleave the transversal CNOT operator with code automorphisms. In-block CNOT operators are then achieved by use of an auxiliary code block (see LABEL:supmat:subsec:full_cnot_group_gen). To demonstrate efficient compilation, we develop substantial linear algebra machinery for certain matrix decomposition problems – we expect these methods to be useful for logical gate compilation in many other code families.
A summary of depth bounds for a range of logical CNOT operators of SHYPS codes is available in LABEL:supmat:tab:cnot_depth.
Notation: We adopt the notation discussed in Example VIII.1, where CNOT circuits on $b$ blocks of $k=r^{2}$ logical qubits or $n=n_{r}^{2}$ physical qubits, are described by invertible matrices in $GL_{br^{2}}(2)$ or $GL_{bn_{r}^{2}}(2)$, respectively. The collection of all (not necessarily invertible) $n\times n$ binary matrices is denoted by $M_{n}(2)$ and $E_{i,j}\in M_{n}(2)$ denotes the matrix of all zeros except the $(i,j)$ entry equal to 1. The collection of diagonal matrices are denoted $\text{Diag}_{n}(2)\subset M_{n}(2)$.
Lemma IX.1.
Let $g_{1},g_{2}\in GL_{r}(2)$. Then the logical CNOT circuits $\left(\begin{matrix}I&g_{1}\otimes g_{2}\\ 0&I\end{matrix}\right),\left(\begin{matrix}I&0\\ g_{1}\otimes g_{2}&I\end{matrix}\right)\in GL_{2r^{2}}(2)$ on code $SHYPS(r)$ may be implemented by a depth-1 physical CNOT circuit.
Proof.
First recall that (like all CSS codes) the transversal $CNOT^{\otimes n_{r}^{2}}$ implements $\overline{CNOT}^{\otimes r^{2}}$ between two code blocks of $SHYPS(r)$. Independently, it follows from Lemmas VIII.6 and VIII.12 that there exist $\sigma_{i}\in S_{n_{r}}$, such that the physical CNOT circuit given by $\left(\begin{matrix}I&0\\ 0&\sigma_{1}\otimes\sigma_{2}\end{matrix}\right)\in GL_{2n_{r}^{2}}(2)$ implements $\overline{\left(\begin{matrix}I&0\\ 0&g_{1}\otimes g_{2}\end{matrix}\right)}\in GL_{2r^{2}}(2).$ Hence
$\left(\begin{matrix}I&0\\ 0&\sigma_{1}^{-1}\otimes\sigma_{2}^{-1}\end{matrix}\right)\cdot\left(\begin{matrix}I&I\\ 0&I\end{matrix}\right)\cdot\left(\begin{matrix}I&0\\ 0&\sigma_{1}\otimes\sigma_{2}\end{matrix}\right)=\left(\begin{matrix}I&\sigma_{1}\otimes\sigma_{2}\\ 0&I\end{matrix}\right),$
implements
$\overline{\left(\begin{matrix}I&0\\ 0&g_{1}^{-1}\otimes g_{2}^{-1}\end{matrix}\right)}\cdot\overline{\left(\begin{matrix}I&I\\ 0&I\end{matrix}\right)}\cdot\overline{\left(\begin{matrix}I&0\\ 0&g_{1}\otimes g_{2}\end{matrix}\right)}=\overline{\left(\begin{matrix}I&g_{1}\otimes g_{2}\\ 0&I\end{matrix}\right)}.$
But recall that $\pi=\sigma_{1}\otimes\sigma_{2}$ is a permutation matrix, and in particular has row and column weights equal to one. Hence
$\left(\begin{matrix}I&\pi\\ 0&I\end{matrix}\right)$
represents the depth-1 physical CNOT circuit $\prod_{i}CNOT_{i,\pi^{-1}(i)+n_{r}^{2}}$. The circuit diagram illustrating this conjugated CNOT operator is given by LABEL:supmat:fig:generalised-transversal-cnot.